OpenStudy (anonymous):

Help please Find the equation of the axis of symmetry and the coordinates of the vertex of the graph of the function y = –2x^2 +6x – 1. A) x = 3; vertex: (3, 35) B) x = –1.5; vertex: (–1.5, – 14.5) C) x = –1.5; vertex: (–1.5, –5.5) D) x = 1.5; vertex: (1.5, 3.5)

OpenStudy (popitree):

from the equation, after simplifying u will get \[(x-\frac{ 3 }{ 2 })^{2} = -\frac{ y }{ 2 }+\frac{ 7 }{ 4 }\]

OpenStudy (popitree):

now can u solve

Nnesha (nnesha):

this question is same just like we did on the previous thred \[\huge\rm \frac{ -b }{ 2a }\] use this formula to find x-coordinate of the vertex

OpenStudy (anonymous):

IN the general quadratic\[y=ax^2 + bx + c\]the axis of symmetry is the vertical lineline\[x=-\frac{ b }{ 2a }\]Yhis also gives the x-coordinate of the vertex. Plug this value of x back into the function to get the y-coordinate of the vertex.

Nnesha (nnesha):

and then replace x by its value into the equation to find y-coordinate axis of symmetry = x coordinate of the vertex

OpenStudy (anonymous):

can we go over this step by step.. still confused So I replace b with -2 and a with 6? @Nnesha

Nnesha (nnesha):

no \[\huge\rm Ax^2+Bx+C=0\] B is middle term A is leading coefficient

OpenStudy (anonymous):

Oh actually the opposite I mean.. -2 -> a and 6 -> b

Nnesha (nnesha):

yes!!

Nnesha (nnesha):

substitute b and a by their values into the formula

OpenStudy (anonymous):

-(6)/2(-2)

Nnesha (nnesha):

yes!

OpenStudy (anonymous):

so now we divide by -2?

Nnesha (nnesha):

there is 2(-2) you can multiply them first or reduce the fraction

OpenStudy (anonymous):

-(6)/-4

Nnesha (nnesha):

yes right :=)

Nnesha (nnesha):

can you reduce the fraction -6/-4 = ?

OpenStudy (anonymous):

so now we simplify

OpenStudy (anonymous):

yes divide by 2 or -2

Nnesha (nnesha):

okay so -6/-4 = ?

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

and then 3/2 = 1.5 right?

Nnesha (nnesha):

yes right

Nnesha (nnesha):

now find y-coordinate of the vertex

OpenStudy (anonymous):

y = -2(1.5)^2 + 6(1.5) - 1?

Nnesha (nnesha):

yes

OpenStudy (anonymous):

= 3.5

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

D) x = 1.5; vertex: (1.5, 3.5)

Nnesha (nnesha):

yes right!

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

my pleasure